# Utilizing Large Language Models to Optimize the Detection and Explainability of Phishing Websites

### v1

@@ 5.1. Evaluating Problem-space evasions
Apruzzese (Yuan et al., 2023) and Yuan et al. (Yuan et al., 2024a) have conceptualized evasive attacks that are relevant and generalizable for creating phishing samples capable of evading the four most popular ML-PWDs. They identified 57 features that attackers commonly use to bypass ML-PWD systems. Building on this framework, Montaruli et al. (Montaruli et al., 2023) developed a novel set of 16 evasive attacks focusing on the problem space of adversarial attacks. These attacks involve directly modifying the HTML content using fine-grained manipulations, allowing for changes to the HTML code of phishing webpages without compromising their maliciousness or visual appearance. This approach successfully preserves both the functionality and appearance of the websites. Some examples of these attacks include InjectIntElem, which involves injecting a specified number of (hidden) internal HTML elements (e.g., ¡a¿ tags with internal links) into the body of the webpage and ObfuscateJS, where the JavaScript code within ¡script¿ elements of a webpage is obfuscated by encoding them into Base64 and then inserts a new script to decode and execute the original script. A full description of these attacks is provided in https://tinyurl.com/vcjer5my). For brevity, we do not go into comprehensive details about which feature (based on Apruzzese (Yuan et al., 2023) and Yuan et al.(Yuan et al., 2024a)) these attacks target.
Since different phishing websites can be compromised by different attacks, Montaruli et al also developed a query-efficient black-box optimization algorithm based on WAF-A-MoLE (Demetrio et al., 2020). This algorithm employs an iterative methodology that involves successive rounds of mutation to modify the original malicious sample, aiming to reduce the confidence score provided by ML-PWD. Attacks are carried out as Single Round (SR) or Multi Round (MR) manipulations. SR manipulations, such as UpdateHiddenDivs, UpdateHiddenButtons, UpdateHiddenInputs, and UpdateTitle, are applied once to achieve their goal of hiding or modifying specific HTML elements and do not require further changes to evade detection. In contrast, MR manipulations, including InjectIntElem and InjectExtElem, require multiple sequential applications to progressively adjust the internal-to-external link ratios and other features to effectively evade detection. The optimizer begins by initializing the best adversarial example and score with the initial phishing webpage and its corresponding score. It then sequentially applies the SR manipulations, updating the best example and score whenever a new manipulation achieves a lower score. Afterward, the optimizer enters the MR manipulations loop, which includes a specified number of mutation rounds. In each mutation round, the algorithm generates new candidate adversarial phishing webpages from the current best example by applying one MR manipulation to each candidate. The candidate with the lowest confidence score is selected, and if this score is lower than the current best, it becomes the new best adversarial example. The number of mutation rounds, $R$, given the maximum query budget $Q$, is determined by the formula $R=\frac{Q-\#SR}{\#MR}$, where #$\#SR$ and #$\#MR$ represent the number of SR and MR manipulations, respectively. One attack InjectFakeFavicon (an SR manipulation) is not relevant to PhishLang since our parser does not look for favicons in <head> tags.
@@
By analyzing these differences, we can identify which manipulations had the most significant impact on reducing the model’s confidence, thereby determining the primary attack that achieved the desired adversarial effect. If two or more attacks demonstrate similar adversarial advantages (within one standard deviation of the mean adversarial advantage across all samples), we designate these as primary attacks [Scenario 1]. Additionally, an attack following a primary attack may negatively impact the model’s performance, but the preceding primary attack might have lessened its effectiveness. Therefore, once a primary attack is identified, we remove it from the sample and rerun the optimization algorithm. If another attack then exhibits an adversarial advantage similar to the previously recognized primary attack (within one standard deviation of the adversarial advantage across all samples), it is also designated as a primary attack [Scenario 2].
In our dataset, we noticed that only 11.3% of the samples met [Criteria 1] and 4.5% (samples met [Criteria 2], indicating that for most phishing websites, only one attack had a significant impact in adversarially enhancing it.
Table 6 presents the performance of each adversarial attack on our sample, detailing the number of incorrect predictions (false negatives) and the corresponding median adversarial advantages. We observe that all problem-space-facing attacks were able to decrease the efficiency of our vanilla model initially, which was then significantly improved by applying the patches. For example, the ObfuscateJS (A8) attack had a median adversarial advantage of 0.407, which dropped to 0.052 after applying adversarial patches, and the incorrect prediction rate reduced from 21.28% to 2.07%. Similarly, the UpdateHiddenButtons (A12) attack showed a median advantage of 0.276 initially, which was reduced to 0.040 post-patch application, with incorrect predictions decreasing from 18.35% to 1.84%.
@@
InjectIntElem (A1) 0.219 (0) 8.02% (0)
InjectIntElemFoot (A2) 0.145 (0) 11.63% (0)
InjectIntLinkElem (A3) 0.267 (0) 4.12% (0)
InjectExtElem (A4) 0.312 (0) 7.89% (0)
InjectExtElemFoot (A5) 0.189 (0) 14.53% (0)
UpdateForm (A6) 0.334 (0) 5.79% (0)
ObfuscateExtLinks (A7) 0.378 (0.047) 19.01% (1.68%)
ObfuscateJS (A8) 0.407 (0.052) 21.28% (2.07%)
@@
UpdateHiddenDivs (A11) 0.198 (0) 8.33% (0)
UpdateHiddenButtons (A12) 0.276 (0.040) 18.35% (1.84%)
UpdateHiddenInputs (A13) 0.322 (0.035) 14.21% (1.75%)
UpdateTitle (A14) 0.154 (0.018) 10.77% (0.98%)
UpdateIFrames (A15) 0.165 (0) 6.15% (0)
@@ 5.2. Mitigating Adversarial Attacks
Adversarial retraining Attack ObfuscateJS (A8) obfuscates JS within ¡script¿ elements to evade the detection features such as HTML_popUP, HTML_rightClick, and HTML_statBar. This technique involves extracting the original JavaScript, encoding it in Base64, and replacing the original <script> content with new code that decodes and executes the obfuscated JavaScript. To counter this attack, we developed two patches: Patch 4.1 converts Base64 (or any other non-UTF-8 encodings) to UTF-8 before parsing, and Patch 4.2 retrains the model with adversarial samples where A7 was the primary attack. Although these patches did not completely nullify the attack, they significantly reduced the median adversarial advantage to 0.047 (from 0.378) and the false negative rate to 1.68% (from 19.01%). Attack UpdateHiddenButtons (A12) obfuscates all disabled button elements by removing the disabled attribute and inserting a new <script> element that re-applies this attribute during rendering using the setAttribute() DOM method. This approach ensures that both the rendering and original behavior of the buttons are preserved. While a potential countermeasure could involve ignoring scripts with the disabled attribute, this could break the user experience. For example, in a web form where the “Submit” button is initially disabled and only enabled after all required fields are filled out, ignoring such scripts could prevent proper functionality. Therefore, to address this attack, we retrained our model with adversarial samples where A12 was the primary attack (Patch 5). This reduced the median adversarial advantage to 0.40 (from 0.276) and the false negative rate to 1.84% (from 18.35%). Attack UpdateHiddenInputs (A13) evades detection of hidden and disabled input elements by changing the type attribute from “hidden” to “text” and adding the hidden attribute. We adversarially trained with samples that had this attack as the primary focus (Patch 6), which reduced the median adversarial advantage to 0.035 (from 0.322) and the false negative rate to 1.75% (from 14.21%). For Attack UpdateTitle (A14), we use a variant of Patch 6 that prevents JS from executing when reading the <title> tag.
@@ 5.3. Advesarial impact on phishing categories
Regular attacks (5,734) 0.207 642 0.024 35
Behaviour based JS (198) 0.418 71 0.239 28
Clickjacking (485) 0.225 88 0.027 13
DOM Manipulations (101) 0.373 43 0.018 2
Text encoding (207) 0.310 41 0.057 7
@@
5.4. Re-evaluating model robustness
We re-evaluated our adversarially patched model on the metadata of the 42.7M domains collected from Certstream that was collected in Section 4.1. We see a noticeable increase in the websites flagged as phishing - 27,574, a 6.88% increase in detection from when the initial model was run. Breaking down these numbers per category: The model detected 18,549 regular (6.6% increase), 3,391 JavaScript evasion (7.3% increase), 3,583 Clickjacking (7.0% increase), 1,147 DOM (8.5% increase), and 904 Text encoding attacks (8.3% increase). However, it is more important to evaluate the False positive and False negative rates of these samples. We pseudo-randomly selected the same number of URLs (2.5k) from the detected set, with the same per category distribution, with 1,632 regular attacks, 280 BJEs, 313 clickjacking, 94 DOM Manipulations, and 78 text encoding. Similar to the evaluation of the initial model, we also randomly selected 2.5k benign websites. Table 8 illustrates the performance of the model against different categories of attacks. We see a noticeable boost across all attack categories, for example, for Behavior-based JS evasion attacks, the model now achieves an accuracy of 94.8% (an increase of 1.9%), whereas for Text Encoding attacks, the accuracy increased to 95.7% (an increase of 3.2%). This indicates that the boost in detection for attack categories transferred from the test set (as seen in the previous section) to real-world websites.
@@ 5.4. Re-evaluating model robustness
6. Evaluating misdetections
Figure 3. Example of an Instagram attack where the seemingly benign landing page (which was missed by PhishLang) led to a phishing page.
In this section, we evaluate the misdetections (False Positives and False Negatives) of PhishLang for both the vanilla model and after adversarial retraining on the 2.5K websites detected as phishing by PhishLang. We also analyzed the reasons for false positives and false negatives, observed whether adversarial improvements introduced additional false negatives, and monitored changes in the false positive and false negative rates for evasive attack categories post-adversarial training.
6.1. False Negatives
Out of 261 undetected websites by PhishLang, 82 had CAPTCHAs and 36 had QR codes, both common phishing evasion tactics (Vidas et al., 2013; Kang and Xiang, 2010). PhishLang successfully detected all phishing websites when QR code URLs were scanned and most CAPTCHA-protected sites, except two inactive ones. For QR codes, we used the pyzbar(Hudson, [n. d.]) library to automate scanning. However, solving CAPTCHAs programmatically can violate many websites’ terms of service (Sivakorn et al., 2016), so PhishLang detects phishing post-CAPTCHA. We also found 25 sites that mimicked legitimate brands but initially only had input fields without malicious behavior. PhishLang’s confidence was low on these landing pages, resulting in false negatives, but it detected phishing on the subsequent pages. Figure 3 shows an Instagram phishing page where PhishLang’s confidence increased from 0.41 on the first page to 0.98 on the second. LIME analysis revealed PhishLang was influenced by brand-related text and form fields but lacked other phishing indicators initially. To address this, our framework now inputs placeholder data on benign-marked sites, submits it, and reassesses the destination, which helped detect all previously missed sites. Additionally, 118 undetected websites were in non-English languages: Japanese (57), Spanish (36), German (18), and Italian (7). Figure 4 shows an example in Japanese. Using langdetect, we found our training data was predominantly in English (n=18,683) with minimal coverage of these languages (Japanese=500, German=660, Spanish=403, and Italian=85). This lack of diversity likely caused misclassifications. While most phishing websites are in English, non-English phishing scams are emerging (Darktrace, 2024).
Post-Adversarial Improvements: Implementing the patches from Section 5 significantly improved performance. Only 2 out of 36 QR-code attacks, 5 out of 25 multi-page attacks, and 7 true-positive detections were misclassified as benign, despite a median confidence score of 0.44.
Figure 4. Example of a non-English (Japanese) false negative sample
6.2. False Positives
Upon evaluating the 112 false positives, we identified 42 malformed websites with poorly structured text and input fields soliciting sensitive information. For example, Fig  5 shows a website with invalid characters and poorly designed input fields for personal information like name, email, and phone number, lacking clear instructions or context. These issues likely caused the site to resemble phishing samples, leading to incorrect classification by PhishLang. Additionally, 23 websites featured multiple empty links and malformed meta tags with several redirections. Addressing these false positives without compromising PhishLang’s detection features is challenging. We also discovered 14 websites with simple login fields that did not transmit data and contained no other information. These sites were classified as false positives because they neither imitated any organization nor had a deceptive motive, despite fitting traditional phishing profiles. Our crawler detects websites as soon as they appear on Certstream, so attackers could later add more details to these sites. We found 21 websites with domain parking pages, likely detected due to their lack of substantial content, presence of multiple advertising links, and frequent redirections. Figure 6 shows one such example. To assess PhishLang’s robustness against parked pages, we used keywords typical of parked pages to select 1,000 parked websites from our total sample of 42.7 million domains. PhishLang flagged only 5 of them as phishing (median confidence of 0.27). Lastly, we noted 12 websites lacking noticeable phishing artifacts, with low confidence scores (Median=0.52). The small number of such samples might not significantly impact the training outcome if included in the dataset. We plan to retrain the model after gathering enough samples matching this criterion.
Post-Adversarial Improvements: Implementing adversarial improvements discussed in Section 5, PhishLang correctly identified 28 of the websites with poor structural design as benign. Out of the 2.5k manually observed true positives, no new samples were changed from phishing to benign, increasing the true positive rate of our model (on the initial dataset) to 96.6%.
Figure 5. Example of a poorly designed (benign)website that was detected as a false positive
Figure 6. Example of a parked website that was detected by PhishLang
@@ 7. Explainable blocklisting
Users often overlook or dismiss browser warnings (Akhawe and Felt, 2013), and even when they do receive antiphishing alerts, these notifications typically lack substantive information beyond simply flagging a website as suspicious (Voronkov et al., 2017). As a result, users are left in the dark, unable to understand the specific reasons behind the warning or learn how to identify potential threats in the future. Our goal is to utilize PhishLang to build a system that can not only identify the features that are suspicious in a phishing website detection but also provide contextual information regarding why they are suspicious. Work by Desolda et al. (Desolda et al., 2023) has explored how explainable blocklisting in phishing websites can aid users. However, not only is their system restricted to limited URL features, which with the presence of newer evasive techniques, is not a good indicator of phishing attacks, but also their system is reliant on generating static messages based on these features, which does not provide users with much context. Recognizing the limitations of existing warning systems, our method utilizes the capabilities of GPT-3.5 Turbo, in conjunction with our DistilBERT-trained model to provide users with comprehensive contextual information when they encounter a phishing alert. Rather than merely informing users that a website has been identified as malicious, our approach analyzes the suspicious features of the flagged site and pinpoints their locations on the website.
@@ 7.1. Generating explainable features
When our DistilBERT model detects a website as phishing, it utilizes LIME to pinpoint the most influential features in its prediction, in an approach similar to the one illustrated in Section 7.3. We consider identifying the top three tags that most significantly influenced the model’s decision and select the tags within the website’s parsed source code that contained those phrases. We chose three tags since we wanted the model to generate explanations for at least three features. However, this value can be increased or decreased as required. These tags, together with the entire parsed source code and the site’s URL, are integrated into a custom prompt that we provide to GPT-3.5T using the OpenAI API. Our framework is illustrated in Figure 8. The purpose of this prompt is to enable the language model to create a dialogue that elaborates on the contextual significance of these suspicious features and their exact locations on the website. Figure 7 illustrates a phishing page designed to imitate Facebook’s Privacy Center that was detected by PhishLang, and the subsequent detailed warning was generated. The system identifies several key indicators of phishing on the page. Firstly, it notes the use of language designed to alarm and hurry the user, a common tactic in phishing schemes aimed at prompting users to disclose personal information impulsively (Alsharnouby et al., 2015). Secondly, the tool highlights that the links on the sidebar are non-functional, another tactic that is used by attackers to make the website look legitimate (Li et al., 2019). Thirdly, PhishLang flags noticeable grammatical mistakes such as ”We have suspended your page,” which are not typical of professional communication from a well-established company like Facebook. For each feature, the system also pinpoints where it appears on the webpage, and also includes a screenshot of the website, enabling users to visually match the suspicious elements with the explanations provided.

### v2

@@ 5.1. Evaluating Problem-space evasions
Apruzzese (Yuan et al., 2023) and Yuan et al. (Yuan et al., 2024a) have conceptualized evasive attacks that are relevant and generalizable for creating phishing samples capable of evading the four most popular ML-PWDs. They identified 57 features that attackers commonly use to bypass ML-PWD systems. Building on this framework, Montaruli et al. (Montaruli et al., 2023) developed a novel set of 16 evasive attacks focusing on the problem space of adversarial attacks. These attacks involve directly modifying the HTML content using fine-grained manipulations, allowing for changes to the HTML code of phishing webpages without compromising their maliciousness or visual appearance. This approach successfully preserves both the functionality and appearance of the websites. Some examples of these attacks include InjectIntElem, which involves injecting a specified number of (hidden) internal HTML elements (e.g., ¡a¿ tags with internal links) into the body of the webpage and ObfuscateJS, where the JavaScript code within ¡script¿ elements of a webpage is obfuscated by encoding them into Base64 and then inserts a new script to decode and execute the original script. A full description of these attacks is provided in the author’s paper. For brevity, we do not go into comprehensive details about which feature (based on Apruzzese (Yuan et al., 2023) and Yuan et al.(Yuan et al., 2024a)) these attacks target.
@@
By analyzing these differences, we can identify which manipulations had the most significant impact on reducing the model’s confidence, thereby determining the primary attack that achieved the desired adversarial effect. If two or more attacks demonstrate similar adversarial advantages (within one standard deviation of the mean adversarial advantage across all samples), we designate these as primary attacks [Scenario 1]. Additionally, an attack following a primary attack may negatively impact the model’s performance, but the preceding primary attack might have lessened its effectiveness. Therefore, once a primary attack is identified, we remove it from the sample and rerun the optimization algorithm. If another attack then exhibits an adversarial advantage similar to the previously recognized primary attack (within one standard deviation of the adversarial advantage across all samples), it is also designated as a primary attack [Scenario 2]. In our dataset, we noticed that only 11.3% of the samples met [Criteria 1] and 4.5% (samples met [Criteria 2], indicating that for most phishing websites, only one attack had a significant impact in adversarially enhancing it. Table 5 presents the performance of each adversarial attack on PhishLang, detailing the number of incorrect predictions (false negatives) and the corresponding median adversarial advantages. We observe that all problem-space-facing attacks were able to decrease the efficiency of our vanilla model initially, which was then significantly improved by applying the patches. For example, the ObfuscateJS (A8) attack had a median adversarial advantage of 0.348, which dropped to 0.028 after applying adversarial patches, and the incorrect prediction rate reduced from 20.92% to 1.51%. Similarly, the UpdateHiddenButtons (A12) attack showed a median advantage of 0.241 initially, which was reduced to 0.025 post-patch application, with incorrect predictions decreasing from 17.20% to 1.07%.
@@
InjectIntElem (A1) 0.204 (0) 7.99% (0)
InjectIntElemFoot (A2) 0.142 (0) 11.24% (0)
InjectIntLinkElem (A3) 0.231 (0) 4.43% (0)
InjectExtElem (A4) 0.285 (0) 7.02% (0)
InjectExtElemFoot (A5) 0.194 (0) 15.01% (0)
UpdateForm (A6) 0.271 (0) 4.90% (0)
ObfuscateExtLinks (A7) 0.314 (0.039) 19.01% (1.68%)
ObfuscateJS (A8) 0.348 (0.028) 20.92% (1.51%)
@@
UpdateHiddenDivs (A11) 0.173 (0) 8.01% (0)
UpdateHiddenButtons (A12) 0.241 (0.025) 17.20% (1.07%)
UpdateHiddenInputs (A13) 0.403 (0.027) 13.59% (1.18%)
UpdateTitle (A14) 0.107 (0.005) 11.27% (0.63%)
UpdateIFrames (A15) 0.111 (0) 5.42% (0)
@@ 5.2. Mitigating Adversarial Attacks
Adversarial retraining Attack ObfuscateJS (A8) obfuscates JavaScript within <script> elements to evade detection features such as HTML_popUP, HTML_rightClick, and HTML_statBar. This technique involves extracting the original JavaScript, encoding it in Base64, and replacing the original <script> content with new code that decodes and executes the obfuscated JavaScript. To counter this attack, we developed two patches: Patch 4.1 converts Base64 (or any other non-UTF-8 encodings) to UTF-8 before parsing, and Patch 4.2 retrains the model with adversarial samples where A7 was the primary attack. Although these patches did not completely nullify the attack, they significantly reduced the median adversarial advantage to 0.039 (from 0.314) and the false negative rate to 1.68% (from 19.01%). Attack UpdateHiddenButtons (A12) obfuscates all disabled button elements by removing the disabled attribute and inserting a new <script> element that re-applies this attribute during rendering using the setAttribute() DOM method. This approach ensures that both the rendering and original behavior of the buttons are preserved. While a potential countermeasure could involve ignoring scripts with the disabled attribute, this could break the user experience. For example, in a web form where the “Submit” button is initially disabled and only enabled after all required fields are filled out, ignoring such scripts could prevent proper functionality. Therefore, to address this attack, we retrained our model with adversarial samples where A12 was the primary attack (Patch 5). This reduced the median adversarial advantage to 0.p025 (from 0.241) and the false negative rate to 1.07% (from 17.20%). Attack UpdateHiddenInputs (A13) evades detection of hidden and disabled input elements by changing the type attribute from “hidden” to “text” and adding the hidden attribute. We adversarially trained with samples that had this attack as the primary focus (Patch 6), which reduced the median adversarial advantage to 0.0.27 (from 0.403) and the false negative rate to 1.18% (from 13.59%). For Attack UpdateTitle (A14), we use a variant of Patch 6 that prevents JS from executing when reading the <title> tag.
@@ 5.3. Advesarial impact on phishing categories
Regular attacks (5,734) 0.212 648 0.030 37
Behaviour based JS (198) 0.425 73 0.246 29
Clickjacking (485) 0.230 91 0.033 14
DOM Manipulations (101) 0.380 45 0.021 3
Text encoding (207) 0.316 42 0.063 8
@@
We also re-evaluated our adversarially patched model on the metadata of the 42.7M domains collected from Certstream that was collected in Section 4.1. we find that 1,778 more websites were flagged as phishing, a noticeable increase of 6.88% compared to when the initial model was run (25,796 URLs to 27,574 URLs). Breaking down these numbers per category: The model detected 18,549 regular (6.6% increase), 3,391 JavaScript evasion (7.3% increase), 3,583 Clickjacking (7.0% increase), 1,147 DOM (8.5% increase), and 904 Text encoding attacks (8.3% increase). However, it is more important to evaluate the False positive and False negative rates of these samples. We pseudo-randomly selected the same number of URLs (2.5k) from the detected set, with the same per category distribution, with 1,632 regular attacks, 280 BJEs, 313 clickjacking, 94 DOM Manipulations, and 78 text encoding. Similar to the evaluation of the initial model, we also randomly selected 2.5k benign websites. Table 7 illustrates the performance of the model against different categories of attacks. We see a noticeable boost across all attack categories, for example, for Behavior-based JS evasion attacks, the model now achieves an accuracy of 94.8% (an increase of 1.9%), whereas for Text Encoding attacks, the accuracy increased to 95.7% (an increase of 3.2%). This indicates that the boost in detection for attack categories transferred from the test set (as seen in the previous section) to real-world websites.
@@ 6. Explainable blocklisting
Security systems often rely on warning dialogs to convey critical information, particularly when blocking malicious content or changes in system configurations that could impact usability (Raja et al., 2011). However, frequent and generic warnings offer little insight into actual threats, leading to user desensitization and ignored alerts (Voronkov et al., 2017; Akhawe and Felt, 2013; Krol et al., 2012). This problem is prevalent in antiphishing systems integrated into web browsers (saf, 2020; mca, 2022), which typically indicate a dangerous website without explaining the specific reasons for flagging it. Evidence suggests that providing contextual information about threats can significantly enhance user understanding and trust (Desolda et al., 2023; Nguyen et al., 2021), yet this approach is rarely adopted in commercial antiphishing tools, leaving users uninformed. Our goal is to leverage PhishLang to develop a system that not only identifies suspicious features in phishing website detection but also provides contextual explanations for why these features are considered suspicious. While Desolda et al. (Desolda et al., 2023) have explored how explainable blocklisting can aid users in recognizing phishing websites, their system is limited to analyzing a narrow set of URL features. Given the emergence of new evasive techniques, these features alone may not effectively indicate phishing attacks. Additionally, their approach relies on generating static messages based on these features, offering users limited context. Recognizing the shortcomings of current warning systems, our method utilizes the capabilities of GPT-3.5 Turbo in conjunction with our MobileBERT-trained model to deliver comprehensive contextual information when users encounter phishing alerts. Instead of simply notifying users that a website is malicious, our approach analyzes the specific suspicious features of the flagged site and highlights their locations, providing users with a deeper understanding of the threat.
@@ 6.1. Generating explainable features
When our MobileBERT model detects a website as phishing, it utilizes LIME to pinpoint the most influential features in its prediction, in an approach similar to the one illustrated in Section 6.3. We consider identifying the top three tags that most significantly influenced the model’s decision and select the tags within the website’s parsed source code that contained those phrases. We chose three tags since we wanted the model to generate explanations for at least three features. However, this value can be increased or decreased as required. These tags, together with the entire parsed source code and the site’s URL, are integrated into a custom prompt that we provide to GPT-3.5T using the OpenAI API. Our framework is illustrated in Figure 3. The purpose of this prompt is to enable the language model to create a dialogue that elaborates on the contextual significance of these suspicious features and their exact locations on the website. Figure 2 illustrates a phishing page designed to imitate Facebook’s Privacy Center that was detected by PhishLang, and the subsequent detailed warning was generated. The system identifies several key indicators of phishing on the page. Firstly, it notes the use of language designed to alarm/rush the user, a common tactic in phishing schemes aimed at prompting users to disclose information impulsively (Alsharnouby et al., 2015). Secondly, the tool highlights that the links on the sidebar are non-functional, another tactic that is used by attackers to make the website look legitimate (Li et al., 2019). Thirdly, PhishLang flags noticeable grammatical mistakes which are not typical of professional communication from a well-established company like Facebook. For each feature, the system pinpoints where it appears on the webpage and includes a screenshot, enabling users to visually match the suspicious elements with the explanations provided.